# Some Aspects of Three-Quark Potentials (Part II)

### v2

@@ Appendix B A static Nambu-Goto string with fixed endpoints
[…]
@@
In general, the tangent angle $\alpha$ may be positive or negative. For $\alpha>0$, the function $r(x)$ describing a string profile is monotonically increasing on the interval $[0,\ell]$. Conversely, for $\alpha<0$, the situation is more intricate. The function $r(x)$ is increasing on $[0,x_{\text{\tiny 0}}]$ and decreasing on $[x_{\text{\tiny 0}},\ell]$, reaching a local maximum at $x=x_{\text{\tiny 0}}$. Both cases are depicted in Figure 13.
@@
that implies $\mathbf{e}_{r}=-T^{-1}\delta S/\delta r_{v}=-\mathsf{g}w(r_{v})\sin\alpha$. Similarly, in static gauge $\xi^{1}=t$ and $\xi^{2}=r$, the boundary term provides the $x$-component of $\mathbf{e}$: $\mathbf{e}_{x}=-\mathsf{g}w(r_{v})\cos\alpha$ [^21] Alternatively, one can choose a gauge in which the boundary terms provide both components a3Q2016 . Putting both components together, we have
@@ 1 The case $\alpha\geq 0$
$E_{R}=\frac{S_{R}}{T}=\mathsf{g}\sqrt{\frac{\mathsf{s}}{v}}\int^{1}_{\sqrt{\tfrac{\mathsf{s}}{v}}\epsilon}\,\frac{du}{u^{2}}\,\text{e}^{vu^{2}}\Bigl{[}1-\cos^{2}{}\alpha\,u^{4}\,\text{e}^{2v(1-u^{2})}\Bigr{]}^{-\frac{1}{2}}\,.$ (B.9)
@@
$E=\mathsf{g}\sqrt{\frac{\mathsf{s}}{v}}\int^{1}_{0}\,\frac{du}{u^{2}}\,\biggl{(}\text{e}^{vu^{2}}\Bigl{(}1-\cos^{2}{}\alpha\,u^{4}\,\text{e}^{2v(1-u^{2})}\Bigr{)}^{-\frac{1}{2}}-1-u^{2}\biggr{)}+c=\mathsf{g}\sqrt{\mathsf{s}}\,{\cal E}^{+}(v,\alpha)+c\,.$ (B.11)
Here the function ${\cal E}^{+}$ is defined in (A.7), and $c$ is a normalization constant.
@@ 2 The case $\alpha\leq 0$
The analysis from the previous subsection generalizes straightforwardly to the case $\alpha\leq 0$. A key point, relevant to all expressions below, is that the string configuration involves two segments: one over the interval $[0,x_{\text{\tiny 0}}]$, where $r(x)$ increases, and another over the interval $[x_{\text{\tiny 0}},\ell]$, where $r(x)$ decreases (see the Figure).
@@
Here $\lambda=\mathsf{s}r_{\text{\tiny 0}}^{2}$, and the function ${\cal L}^{-}$ is given by (A.3). It is important to note that $\lambda$, $v$, and $\alpha$ are not independent. From the first integral it follows that
@@ Appendix C Some details on the quark-antiquark potential
Here $\lambda=\mathsf{s}r_{\text{\tiny 0}}^{2}$ is a parameter varying from 00 to $1$, and $c$ is the same normalization constant as in Appendix B. The functions ${\cal L}^{+}$ and ${\cal E}^{+}$ are defined in Appendix A.
The $-\ell$ behavior of $E_{\text{\tiny Q}\bar{\text{\tiny Q}}}$ is given by
@@
In contrast, the $-\ell$ behavior is given by
@@
The factor $I_{0}$ is defined in Appendix A. It is important to note that the coefficients $\boldsymbol{\sigma}_{\text{\tiny Q}\bar{\text{\tiny Q}}}$ and $\sigma$ differ, with their ratio given numerically by $\boldsymbol{\sigma}_{\text{\tiny Q}\bar{\text{\tiny Q}}}/\sigma=0.805$. Furthermore, the difference between the constant terms in the small- and $-\ell$ expansions is $2c-C_{\text{\tiny Q}\bar{\text{\tiny Q}}}=175\,\text{MeV}$, indicating that the constant term in the $-\ell$ expansion is larger.
@@ D.1 The small $\ell$ limit
To gain some insights, we begin by examining the behavior of string configurations at small $\ell$. As shown in Appendix E of a3Q2016 , in the diquark limit (small $\theta$), the potential is described by configuration I. This remains valid as long as the parameter $v$ is less than ${\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}$, where ${\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}$ is the solution to the equation $\alpha_{1}(v)=0$. We want to find the value of $\theta$ such that the equation has the trivial solution ${\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}=0$. In other words, the interval $[0,{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}]$ shrinks to a point, rendering configuration I irrelevant. Substituting $\alpha_{1}=0$ into Eqs.(3.5) and (3.6), we find that
@@ D.2 The range $0<\theta<\theta^{1}$
=$\ell=\begin{cases}\ell^{\,\text{\tiny I}}(v)\,,\,\,&0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny I}}(v)\,,\,\,&0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\\ \end{cases}\quad\text{for}\quad 0<\theta\leq 0.296\pi\,,\,0.298\pi\leq\theta<\theta^{1}\,,$ (D.4)
@@
=$\ell=\begin{cases}\ell^{\,\text{\tiny I}}(v)\,,\,\,&0\,\leq\,v\leq\,{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 1}}\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 1}}\leq v\leq{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 1}}\,,\\ \ell^{\,\text{\tiny I}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 1}}\,\leq v\leq{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 1}}\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 1}}\leq\,v\leq\,\,{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny I}}(v)\,,\,\,&0\,\leq\,\,v\leq\,{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 1}}\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 1}}\leq v\leq{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 1}}\,,\\ E^{\,\text{\tiny I}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 1}}\,\leq v\leq{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 1}}\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 1}}\leq\,v\leq\,\,{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad\,\,\,\text{for}\quad 0.296\pi<\theta<0.298\pi\,.$ (D.5)
@@ D.3 The range $\theta^{1}\leq\theta\leq\theta^{3}$
=$\ell=\begin{cases}\ell^{\,\text{\tiny II}}(v)\,,\,\,&0\,\leq\,\,v\leq{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\,\,,\\ \ell^{\,\text{\tiny III}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\leq\,v\leq{\text{v}}_{\text{\tiny 1}}\,\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny II}}(v)\,,\,\,&0\,\leq\,\,v\leq{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\,\,,\\ E^{\,\text{\tiny III}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\,,\qquad\text{for}\quad 0.350\pi<\theta\leq\theta^{3}\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\leq\,v\leq{\text{v}}_{\text{\tiny 1}}\,\,.\end{cases}$ (D.8)
@@
=$\ell=\begin{cases}\ell^{\,\text{\tiny III}}(v)\,,\,\,&0\,\leq\,\,v\leq{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\,\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\,,\\ \ell^{\,\text{\tiny III}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 3}}\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 3}}\leq\,v\leq{\text{v}}_{\text{\tiny 1}}\,\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny III}}(v)\,,\,\,&0\,\leq\,\,v\leq{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\,\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\,,\\ E^{\,\text{\tiny III}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 3}}\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 3}}\leq\,v\leq{\text{v}}_{\text{\tiny 1}}\,\,,\end{cases}\qquad\text{for}\quad\theta^{3}<\theta\leq 0.353\pi\,,$ (D.9)
@@
=$\ell=\begin{cases}\ell^{\,\text{\tiny III}}(v)\,,\,\,&0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny III}}(v)\,,\,\,&0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad\text{for}\quad 0.353\pi<\theta<0.66638\pi\,.$ (D.10)

### v3

@@ Appendix B A static Nambu-Goto string with fixed endpoints
$S=T\mathsf{g}\int_{0}^{\ell}dx\,w(r)\sqrt{1+(\partial_{x}r)^{2}}\,,\qquad w(r)=\frac{\text{e}^{\mathsf{s}r^{2}}}{r^{2}}\,.$ (B.2)
For convenience, we use the shorthand notation $\mathsf{g}=\frac{R^{2}}{2\pi\alpha^{\prime}}$, $T=\int dt$, and $\partial_{x}r=\frac{\partial r}{\partial x}$. Since the integrand does not depend explicitly on $x$, the equation of motion admits a first integral
@@
In general, the tangent angle $\alpha$ may be positive or negative. For $\alpha>0$, the function $r(x)$ describing a string profile is monotonically increasing on the interval $0,\ell]$. Conversely, for $\alpha<0$, the situation is more intricate. The function $r(x)$ is increasing on $0,x_{\text{\tiny 0}}]$ and decreasing on $x_{\text{\tiny 0}},\ell]$, reaching a local maximum at $x=x_{\text{\tiny 0}}$. Both cases are depicted in Figure 13.
@@
that implies $\mathbf{e}_{r}=-T^{-1}\delta S/\delta r_{v}=-\mathsf{g}w(r_{v})\sin\alpha$. Similarly, in static gauge $\xi^{1}=t$ and $\xi^{2}=r$, the boundary term provides the $x$-component of $\mathbf{e}$: $\mathbf{e}_{x}=-\mathsf{g}w(r_{v})\cos\alpha$. [^21] Alternatively, one can choose a gauge in which the boundary terms provide both components a3Q2016 . Putting both components together, we have
@@ 1 The case $\alpha\geq 0$
$\ell=\cos\alpha\sqrt{\frac{v}{\mathsf{s}}}\int^{1}_{0}du\,u^{2}\,\text{e}^{v(1-u^{2})}\Bigl(1-\cos^{2}{}\hskip-2.84526pt\alpha\,u^{4}\text{e}^{2v(1-u^{2})}\Bigr)^{-\frac{1}{2}}=\frac{1}{\sqrt{\mathsf{s}}}{\cal L}^{+}(\alpha,v)\,$, (B.8)
@@
$E_{R}=\frac{S_{R}}{T}=\mathsf{g}\sqrt{\frac{\mathsf{s}}{v}}\int^{1}_{\sqrt{\tfrac{\mathsf{s}}{v}}\epsilon}\,\frac{du}{u^{2}}\,\text{e}^{vu^{2}}\Bigl[1-\cos^{2}{}\hskip-2.84526pt\alpha\,u^{4}\,\text{e}^{2v(1-u^{2})}\Bigr]^{-\frac{1}{2}}\,.$ (B.9)
@@
$E=\mathsf{g}\sqrt{\frac{\mathsf{s}}{v}}\int^{1}_{0}\,\frac{du}{u^{2}}\,\biggl(\text{e}^{vu^{2}}\Bigl(1-\cos^{2}{}\hskip-2.84526pt\alpha\,u^{4}\,\text{e}^{2v(1-u^{2})}\Bigr)^{-\frac{1}{2}}-1-u^{2}\biggr)+c=\mathsf{g}\sqrt{\mathsf{s}}\,{\cal E}^{+}(v,\alpha)+c\,.$ (B.11)
[…]
@@ 2 The case $\alpha\leq 0$
The analysis from the previous subsection generalizes straightforwardly to the case $\alpha\leq 0$. A key point, relevant to all expressions below, is that the string configuration involves two segments: one over the interval $0,x_{\text{\tiny 0}}]$, where $r(x)$ increases, and another over the interval $x_{\text{\tiny 0}},\ell]$, where $r(x)$ decreases (see the Figure).
@@
$\ell=\sqrt{\frac{\lambda}{\mathsf{s}}}\biggl[\int^{1}_{0}du\,u^{2}\,\text{e}^{\lambda(1-u^{2})}\Bigl(1-u^{4}\,\text{e}^{2\lambda(1-u^{2})}\Bigr)^{-\frac{1}{2}}+\int^{1}_{\sqrt{\frac{v}{\lambda}}}du\,u^{2}\,\text{e}^{\lambda(1-u^{2})}\Bigl(1-u^{4}\,\text{e}^{2\lambda(1-u^{2})}\Bigr)^{-\frac{1}{2}}\biggr]=\frac{1}{\sqrt{\mathsf{s}}}{\cal L}^{-}(\lambda,v)\,.$ (B.13)
@@ Appendix C Some details on the quark-antiquark potential
[…]
The small-$\ell$ behavior of $E_{\text{\tiny Q}\bar{\text{\tiny Q}}}$ is given by
@@
In contrast, the large-$\ell$ behavior is given by
@@
The factor $I_{0}$ is defined in Appendix A. It is important to note that the coefficients $\boldsymbol{\sigma}_{\text{\tiny Q}\bar{\text{\tiny Q}}}$ and $\sigma$ differ, with their ratio given numerically by $\boldsymbol{\sigma}_{\text{\tiny Q}\bar{\text{\tiny Q}}}/\sigma=0.805$. Furthermore, the difference between the constant terms in the small- and large-$\ell$ expansions is $2c-C_{\text{\tiny Q}\bar{\text{\tiny Q}}}=175\,\text{MeV}$, indicating that the constant term in the small-$\ell$ expansion is larger.
@@ D.1 The small $\ell$ limit
To gain some insights, we begin by examining the behavior of string configurations at small $\ell$. As shown in Appendix E of a3Q2016 , in the diquark limit (small $\theta$), the potential is described by configuration I. This remains valid as long as the parameter $v$ is less than ${\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}$, where ${\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}$ is the solution to the equation $\alpha_{1}(v)=0$. We want to find the value of $\theta$ such that the equation has the trivial solution ${\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}=0$. In other words, the interval $0,{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}]$ shrinks to a point, rendering configuration I irrelevant. Substituting $\alpha_{1}=0$ into Eqs.(3.5) and (3.6), we find that
@@ D.2 The range $0<\theta<\theta^{1}$
$\ell=\begin{cases}\ell^{\,\text{\tiny I}}(v)\,,\,\,&0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny I}}(v)\,,\,\,&0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 1}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\\ \end{cases}\quad\text{for}\quad 0<\theta\leq 0.296\pi\,,\,0.298\pi\leq\theta<\theta^{1}\,$, (D.4)
@@
$\ell=\begin{cases}\ell^{\,\text{\tiny I}}(v)\,,\,\,&0\,\leq\,v\leq\,{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 1}}\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 1}}\leq v\leq{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 1}}\,,\\ \ell^{\,\text{\tiny I}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 1}}\,\leq v\leq{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 1}}\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 1}}\leq\,v\leq\,\,{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny I}}(v)\,,\,\,&0\,\leq\,\,v\leq\,{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 1}}\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 1}}\leq v\leq{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 1}}\,,\\ E^{\,\text{\tiny I}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 1}}\,\leq v\leq{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 1}}\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 1}}\leq\,v\leq\,\,{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad\,\,\,\text{for}\quad 0.296\pi<\theta<0.298\pi\,.$ (D.5)
@@ D.3 The range $\theta^{1}\leq\theta\leq\theta^{3}$
$\ell=\begin{cases}\ell^{\,\text{\tiny II}}(v)\,,\,\,&0\,\leq\,\,v\leq{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\,\,,\\ \ell^{\,\text{\tiny III}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\leq\,v\leq{\text{v}}_{\text{\tiny 1}}\,\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny II}}(v)\,,\,\,&0\,\leq\,\,v\leq{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\,\,,\\ E^{\,\text{\tiny III}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\,,\qquad\text{for}\quad 0.350\pi<\theta\leq\theta^{3}\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\leq\,v\leq{\text{v}}_{\text{\tiny 1}}\,\,.\end{cases}$ (D.8)
@@
$\ell=\begin{cases}\ell^{\,\text{\tiny III}}(v)\,,\,\,&0\,\leq\,\,v\leq{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\,\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\,,\\ \ell^{\,\text{\tiny III}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 3}}\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 3}}\leq\,v\leq{\text{v}}_{\text{\tiny 1}}\,\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny III}}(v)\,,\,\,&0\,\leq\,\,v\leq{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\,\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 01}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\,,\\ E^{\,\text{\tiny III}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 02}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 3}}\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 03}}^{\text{\tiny 3}}\leq\,v\leq{\text{v}}_{\text{\tiny 1}}\,\,,\end{cases}\qquad\text{for}\quad\theta^{3}<\theta\leq 0.353\pi\,$, (D.9)
@@
$\ell=\begin{cases}\ell^{\,\text{\tiny III}}(v)\,,\,\,&0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\,,\\ \ell^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny III}}(v)\,,\,\,&0\,\leq v\leq{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\,,\\ E^{\,\text{\tiny II}}(v)\,,\,\,&{\text{v}}_{\text{\tiny 0}}^{\text{\tiny 3}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad\text{for}\quad 0.353\pi<\theta<0.66638\pi\,.$ (D.10)